OpenStudy (anonymous):

Eliminate the parameter. x = 3t, y = t + 7

OpenStudy (anonymous):

y = (x/3) + 7 y = 3x + 7 y = 3x - 7 y = (x/3) - 7

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Same process as before. Find a simple substitution based on how the parameter is used in \(x\) and \(y\). Since both \(x\) and \(y\) have \(t\), you can say \[x=3t~\Rightarrow~t=\frac{x}{3}\] Now plug that into \(y\).

OpenStudy (anonymous):

y = x/3 + 7

OpenStudy (anonymous):

Yes